# WiLLM: An Open Wireless LLM Communication System

### v1

@@ Abstract.
The rapid evolution of Large Language Models (LLMs) threatens to overwhelm the existing wireless infrastructure, which calls for architectural innovations and frameworks to support the surging mobile LLM services. This paper proposes WiLLM, the first open-source wireless communication system specifically designed for LLM services, consisting of three progressive innovations. Firstly, we establish a novel paradigm by deploying LLMs in core networks (CNs) with abundant GPUs to enable distributed LLM inference services, strategically positioning LLM inference at the convergence point of the backbone bandwidth and the edges of the cellular network. Secondly, we propose an innovative extension to the conventional network slicing architecture, redefining its design with a “Tree-Branch-Fruit” framework. Its specialized adaptation for LLM services enables telecom operators to monetize these services through slice subscriptions while retaining infrastructure ownership. Finally, to operationalize this architectural vision, we present WiLLM, which systematically addresses critical limitations in existing solutions through novel key capabilities previously unattainable: (1) enhanced slice orchestration, extending existing primary slicing to implement a dual-layer slicing architecture and achieving coordinated multi-UE-multi-slice scheduling mechanisms that provide finer-grained resource allocation for LLM services; (2) universal UE compatibility through application-layer tunneling mechanisms, enabling legacy devices without native slicing capabilities to utilize LLM slice services without hardware upgrades; (3) dual-mode scheduling and cross-layer APIs, supporting flexible deployment from CNs to servers. Built on OpenAirInterface, WiLLM extends this established framework with innovations specific to LLM services, lowering the barrier for adoption by researchers without wireless expertise. Furthermore, we release the first LLM wireless communication dataset with 1,649,996 records and synchronized 58-dimensional metrics, along with two relevant benchmarks. A small but quintessential case study of smart glasses demonstrates practical viability in resource-constrained devices and human-machine interaction scenarios. WiLLM aims to build an open platform and ecosystem for cross-layer optimization and AI-telecom convergence. The code, datasets, hardwares and videos are available at https://openwillm.github.io.

### v2

# WiLLM: an Open Framework for LLM Services over Wireless Systems

@@ Abstract.
Large Language Model (LLM) services fundamentally differ from traditional Deep Neural Network (DNN) applications in wireless networks. We identify three critical distinctions: (1) unlike traditional DNNs with unidirectional data flows, LLM’s multimodal interactions create bidirectional heavy loads with contrasting bottlenecks, requiring direction-aware resource scheduling; (2) while traditional DNNs exhibit fixed computational patterns, LLM’s highly variable inference times interact complexly with network slicing, causing dynamic bottleneck migration; and (3) in contrast to predictable DNN traffic, LLM’s token streams demonstrate unprecedented burstiness and state dependencies. These insights motivate WiLLM, the first open-source framework, implemented as a wireless platform, for LLM service research. Built on OpenAirInterface, WiLLM introduces several technical innovations: dynamic slice compatibility, universal UE compatibility through application-layer tunneling, multi-UE multi-slice scheduling, dual-mode resource allocation, and cross-layer APIs. In addition, WiLLM eliminates the need for specialized wireless expertise, enabling researchers and developers to experiment with LLM services over realistic cellular networks. We demonstrate the platform’s capabilities through a smart glasses case study and provide a comprehensive dataset of 1̃.6 million synchronized measurements. The complete system, dataset, and appendix are available at https://openwillm.github.io.
@@
1. Introduction
The proliferation of Large Language Models across mobile applications creates unprecedented demands on wireless network infrastructure. From augmented reality assistants to real-time translation services, LLM-powered applications require seamless integration of computational inference and wireless communication (Bariah et al., 2024). However, deploying LLM services over cellular networks reveals fundamental mismatches between their operational characteristics and existing network architectures (Boateng et al., 2024; Chaoub and Elkotob, 2025). Modern wireless systems, evolved to handle video streaming and web traffic, struggle with the unique patterns of token generation, multimodal processing, and bidirectional heavy loads that characterize LLM interactions.